Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 395     Entry time: 19 Oct 2023, 23:30
Author: Jayanath Koliyadu 
Group: SPB 
Subject: Shift end 

X-ray delivery

  • 13.6 keV, ~1.7 mJ and up to 352 pulses at 1.1 MHz

Optical laser delivery

  • PP laser, 1 to 202 pulses, stable delivery except for the down time of ~1h

Achievements/Observations

  • user sample data collected for ~1.5h
  • PP laser timing scans done
  • new adapter for the screens holder designed, machined and installed

Issues

  • sample was growing on the adapter, preventing data collection
  • new adapter is too far downstream, partially shadowing sample diffraction on the lower corners of the detector
  • PP laser down for ~1h
ELOG V3.1.4-7c3fd00